### v1

# Scientific Paper Retrieval with LLM-Guided Semantic-Based Ranking

Abstract
Scientific paper retrieval is essential for supporting literature discovery and research. While dense retrieval methods demonstrate effectiveness in general-purpose tasks, they often fail to capture fine-grained scientific concepts that are essential for accurate understanding of scientific queries. Recent studies also use large language models (LLMs) for query understanding; however, these methods often lack grounding in corpus-specific knowledge and may generate unreliable or unfaithful content. To overcome these limitations, we propose SemRank, an effective and efficient paper retrieval framework that combines LLM-guided query understanding with a concept-based semantic index. Each paper is indexed using multi-granular scientific concepts, including general research topics and detailed key phrases. At query time, an LLM identifies core concepts derived from the corpus to explicitly capture the query’s information need. These identified concepts enable precise semantic matching, significantly enhancing retrieval accuracy. Experiments show that SemRank consistently improves the performance of various base retrievers, surpasses strong existing LLM-based baselines, and remains highly efficient.
Scientific Paper Retrieval with LLM-Guided Semantic-Based Ranking
Yunyi Zhang1, Ruozhen Yang1, Siqi Jiao1, SeongKu Kang2, Jiawei Han1 1University of Illinois Urbana-Champaign 2Korea University yzhan238@illinois.edu
1 Introduction
Scientific paper retrieval is a crucial task to facilitate literature discovery and accelerate scientific progress Kang et al. (2024b). Unlike general-purpose information retrieval, scientific paper retrieval is more challenging because queries often involve theme-specific intent and specialized terminology. In addition, acquiring labeled query-passage pairs for supervised fine-tuning is costly and requires domain expertise, making it impractical to continuously annotate more data to adapt the fast-evolving scientific domains.
Recently, dense passage retrieval methods have been widely studied in various ad hoc searches Karpukhin et al. (2020); Izacard et al. (2021). These methods encode the overall semantics of queries and passages into the same vector space and measure relevance using vector similarity. Although being effective in different general domain applications, they still face challenges in scientific paper retrieval.
Specifically, general-purpose semantic representations learned by dense retrievers often fail to capture fine-grained scientific concepts that are crucial for accurately understanding and satisfying a scientific query. For example, the query “Can you point me to studies discussing methods for evaluating text generation models on various dimensions?” involves not only general topics like “natural language generation” and “automatic evaluation” that need to be inferred from the text, but also specific details like “multidimensional evaluation”. A dense retriever, however, only encodes the text in a holistic view, while it lacks the ability and controllability to focus on the scientific concepts which are the core need of the query.
Figure 1: An illustrative example from LitSearch with SPECTER-v2 as base retriever. By capturing the scientific concepts for corpus and query, SemRank substantially improves the ranking results.
With the advancements of large language models (LLM) such as GPT OpenAI (2023) and Claude Anthropic (2024), recent studies also explore how to utilize LLMs in query understanding to help retrieval tasks. For example, HyDE Gao et al. (2023) uses an LLM to generate a hypothetical passage for encoding, and CSQE Lei et al. (2024) prompts an LLM to select a set of relevant sentences to expand the original query. However, these methods still rely on a pre-trained dense retriever to encode overall semantics on the document or sentence levels, lacking the ability to explicitly capture what the query is asking for. In addition, LLMs are not inherently retrieval models. They do not have the vast and dynamic knowledge in the scientific literature necessary to understand scientific queries. Therefore, when used in a zero-shot manner, LLMs cannot identify domain-specific terminology and may generate hallucinated content. Therefore, how to effectively augment LLMs with corpus-based knowledge while capturing specific query information remains a challenge.
To overcome these limitations, we propose to utilize LLMs’ text understanding ability in the scientific retrieval task with the help of a scientific concept-based semantic index. Unlike earlier studies which construct domain-specific semantic index at topic level Tsatsaronis et al. (2015), we build concept-based semantic index at various granularities: from broad research topics such as “natural language generation” to specific terms such as “multidimensional evaluation metrics”. These multi-granular concepts capture the essential content of the paper. Then, we use the semantic index to improve any existing retrievers with the help of LLMs. Specifically, we prompt an LLM to identify a set of core concepts to explicitly represent what the scientific query is asking for. We augment the prompt with candidate concepts derived from the corpus, which helps the LLM to reduce hallucination and ensure the generated content align with the semantic index of corpus. Figure 1 shows an example that, by accurately capturing the multi-granular concepts of the corpus and query, we can improve the scientific paper retrieval results by focusing on the core need of the query.
We propose the SemRank, LLM-Guided Semantic-Based Ranking, a plug-and-play framework for scientific paper retrieval. First, during the indexing time, we build scientific concept-based semantic index for the corpus by identifying a set of research topics and key phrases for each paper. To ensure the topic labels in their canonical forms, we train an auxiliary topic classifier model to identify a set of candidate topics from a large label space [^1] An academic label space is widely available such as ACM CCS and Microsoft Academic Graph Sinha et al. (2015). Then, we prompt an LLM to select the core topic labels and extract key phrases to build the semantic index. Then given a query during retrieval, we first construct a set of candidate concepts from the corpus using a base retriever, from which an LLM can be augmented with corpus-based knowledge and identify a set of core concepts for the query. These concepts, serving as an explicit guidance on what the query is asking for, will be used for concept-level semantic matching to improve the retrieval results.
Figure 2: Overview of the SemRank framework.
Our method can be easily integrated with any dense retriever and improve their retrieval quality without relying on any annotated query data. Experiments show that our method is both effective and efficient. During retrieval time, SemRank only needs one LLM prompting per query, no additional call to the base retriever, and all computation can be done easily on CPUs. Yet it can significantly improve the ranking performance of a wide range of base retrievers and outperforms various LLM-based baselines.
The contributions of this paper are as follows:
  • We propose SemRank, a plug-and-play framework for scientific paper retrieval, which utilizes concept-based semantic index and LLM guidance to explicit capture the information need for scientific queries.
  • We develop a light-weighted method which augments an LLM with semantic index of corpus to accurately identify a set of core scientific concepts for the query, which are then used to improve the retrieval performance through concept-level semantic matching.
  • Through extensive experiments, we show that SemRank consistently improves the base retrievers’ performance and outperforms existing baselines while being more efficient.
2 Problem Formulation
Given a corpus $\mathcal{D}$ of scientific papers, our goal is to retrieve and rank the papers according to their relevance to a given query $q$. Specifically, we will build concept-based semantic index for the corpus by identifying a set of scientific concepts representing the core information of each paper. We assume a scientific topic label space $\mathcal{T}$, which is widely available and normally contains a large number of research topics. In this work, we use Microsoft Academic Graph Sinha et al. (2015), which contains 13,613 Computer Science topics. Then, we utilize LLMs and the semantic index to improve the scientific paper retrieval performance.
3 Methodology
In this section, we will present our SemRank framework. We first introduce the offline semantic index construction module (Sect. 3.1), then we introduce the LLM-guided semantic-based ranking for scientific paper retrieval (Sect. 3.2). Figure 2 shows an overview of SemRank.
3.1 Semantic Index Construction
To capture the core research concepts at various granularities for scientific papers, we propose to build a semantic index of the corpus that contains general research topics and specific key phrases. The broad topics aim to cover the overall themes of the papers that are not explicitly mentioned such as “natural language generation” and “automatic evaluation,” while the key phrases aim to capture the detailed information specific to the paper such as “multidimensional valuation metrics”. Such a well-structured corpus foster flexible query matching at different granularities.
While directly prompting LLMs could be a viable solution to assign each document a set of topics and key phrases, it is hard to ensure the faithfulness of the LLM-generated content. Besides, the research topics are often not explicitly mentioned in the text and LLMs can generate topics at random granularity, which is less controllable and hard to match during retrieval. Therefore, we propose to first fine-tune a multi-label topic classifier for scientific papers with a domain-specific topic label space, from which we can obtain a set of candidate topics for each scientific paper.
Candidate Topic Prediction
We first fine-tune a multi-label text classifier to estimate the likelihood of a paper belonging to a research topic. We use a simple log-bilinear text matching network as our model architecture Zhang et al. (2025). We initialize the paper encoder with a pre-trained scientific domain language model (e.g., SPECTER-v2 Singh et al. (2022)). We also get the topic embeddings using the same pre-trained model and detach them from the encoder. This ensures only the embeddings are updated without back-propagating to the backbone for saving cost. Then, the classifier predicts the probability of document $d_{i}$ belonging to topic $t_{j}\in\mathcal{T}$ with log-bilinear matching:
$p(t_{j}|d_{i})=\sigma(\exp{(\mathbf{t}_{j}^{T}\mathbf{W}\mathbf{d}_{i})}),$
where $\sigma$ is the sigmoid function, $\mathbf{W}$ is a learnable interaction matrix, and $\mathbf{t}_{j}$ and $\mathbf{d}_{i}$ are the encoded topic and document.
We fine-tune the multi-label topic classifier using the binary cross entropy loss. Given the positive topics of paper $d_{i}\in\mathcal{D}$ in the label space $\mathcal{T}_{i}\subset\mathcal{T}$, we train the classifier with:
$\mathcal{L}=-\sum_{d_{i}\in\mathcal{D}}$ $\Big{(}\sum_{t_{j}\in\mathcal{T}_{i}}\log{p(t_{j}|d_{i})}$
$+\alpha\sum_{t_{j}\notin\mathcal{T}_{i}}\log{(1-p(t_{j}|d_{i}))}\Big{)},$
where the coefficient $\alpha$ is a small constant number to counter the imbalanced numbers of positive and negative labels in a large label space.
Core Concepts Identification
With the fine-tuned scientific topic classifier, we first predict a set of candidate topics likely relevant to each paper in the retrieval corpus. Then, we prompt an LLM to perform two tasks: (1) select a set of topics from the candidate list that are not too broad or irrelevant, and (2) extract a set of key phrases from the paper. By providing a candidate topic list, we turn a generation task (of zero-shot prompting) to a selection and extraction task for LLMs, which reduces the chance of hallucination and also ensures the selected topics following a label space. Figure 4 shows our prompt.
In summary, for each scientific paper $d_{i}\in\mathcal{D}$, we identify a set of research topics belonging to a label space, denoted as $T_{i}\in\mathcal{T}$, and a set of key phrases extracted from it, denoted as $P_{i}$.
3.2 LLM-Guided Semantic-Based Retrieval
With the constructed semantic index capturing the core concepts discussed in the corpus, we now present how SemRank leverages it to enhance retrieval by explicitly modeling a query’s information need. Given a base retriever, SemRank first identifies a set of candidate concepts relevant to the query and prompts an LLM to analyze the retrieval context and select the most salient core concepts. These core concepts are then used to refine the initial retrieval results, yielding a concept-aware ranking that better aligns with the query intent.
Candidate Concepts Construction for Query
Directly prompting an LLM to assign a set of scientific concepts to a query is not optimal, because the generated content may not align with the semantic index structure of the corpus and thus make it difficult to match between the query and the papers. Therefore, we propose to first construct a set of candidate concepts for a query from the corpus.
Inspired by the idea of pseudo relevance feedback, given a base retriever $s^{base}$ and its initial ranking results, we collect a set of topics and key phrases that are frequently mentioned in the top-ranked papers. Specifically, for a query $q$, we collect lists of top-k most frequent topics and key phrases mentioned by the top-ranked papers using their pre-constructed semantic index $T_{i}$ and $P_{i}$, and denote the lists as $T^{0}(q)$ and $P^{0}(q)$. These concepts, being frequently mentioned by top-ranked papers, are likely relevant to the initial query.
LLM-Guided Core Concept Identification
With the constructed candidate concepts of the query, we can now prompt an LLM to identify a set of core concepts for the query that can most likely identify its relevant papers. Given the topic and key phrases lists $T^{0}(q)$ and $P^{0}(q)$ as well as the top-k papers in the current ranking $D^{0}(q)$, we instruct an LLM to select a set of concepts from the candidate lists in order to improve the current ranking results. Figure 5 shows our prompt.
Because the candidate concepts are collected from the semantic index, they contribute to two advantages: (1) they serve as a high-level summary of the current retrieval results to help the LLM interpret the query and current results; (2) they provide a high-quality candidate list for the LLM, ensuring the faithfulness and reducing hallucination.
After prompting the LLM, we get a set of core concepts selected from the candidate lists, which we denote as $C(q)$. This set contains the most important concepts relevant to the original query and thus explicitly represents the information need of the query. Also, the concepts are in different granularities, proper for matching at the needed level of details. For example, for the query shown in Fig 1, the LLM identifies general topics like “natural language generation” and “automatic evaluation”, and specific terms like “multidimensional evaluation” and “dialogue response generation”.
Core Concept-Based Ranking
Finally, with the identified set of core concepts of the query, we can use the semantic index of the corpus to re-evaluate the ranking. Specifically, given the core concepts of the query $C(q)$ and the the concepts of a paper $C_{i}=T_{i}\cup P_{i}$, we calculate their similarity with a multi-vector similarity matching score.
$s^{sem}(q,d_{i})=\frac{1}{|C(q)|}\sum_{c\in C(q)}\max_{c^{\prime}\in C_{i}}sim(\mathbf{c},\mathbf{c^{\prime}}),$
where $\mathbf{c}$ denotes the embedding of a concept by a semantic encoder (e.g., SPECTER-v2) and $sim()$ represents the cosine similarity function. This embedding-based soft matching process identifies the most similar concept in a paper for each query’s concept, which accounts for the situation where similar concepts are expressed slightly differently (e.g., "hallucination" and "hallucinated content"). Then, we combine this semantic-based score with the base retriever’s score by z-score normalization (denoted by $z(\cdot)$), based on which we can re-rank the papers to get a new ranked list.
$s(q,d_{i})=z(s^{base}(q,d_{i}))+z(s^{sem}(q,d_{i})).$
Efficiency of SemRank Retrieval
Given the ranking results of a base retriever, our LLM-guided semantic-based ranking process is highly efficient. First, it only requires one LLM call, and the output length is minimal because it is highly-structured with a list of scientific concepts. Second, because all the query’s concepts are selected from the semantic index and all concept embeddings can be pre-computed offline, we only need the cosine similarity computation (i.e., dot product with normalized vectors) during retrieval, which is highly efficient and can be done on CPUs. Besides, the pairwise similarities between concepts can also be pre-computed for the maximal inference efficiency, but it requires substantial amount of storage, so we opt to compute the similarity during retrieval. We conduct a detailed efficiency analysis in Sect. 4.4.
4 Experiments
4.1 Experiment Setup
Table 1: Datasets overview.
Dataset corpus size # test query doc/query
CSFCube 4,207 34 13.32
DORISMAE 8,482 90 19.49
LitSearch 64,183 597 1.07
Datasets
We use three public datasets on scientific paper retrieval: CSFCube Mysore et al. (2021), DORISMAE Wang et al. (2023a), and LitSearch Ajith et al. (2024), including both human-annotated and LLM-generated relevance labels. We use the processed version of CSFCube and DORISMAE released by Kang et al. (2024b) [^2] https://aclanthology.org/attachments/2024.emnlp-main.407.data.zip and LitSearch from its official github [^3] https://github.com/princeton-nlp/LitSearch. Table 1 summarizes the overall statistics of the datasets.
Table 2: Performance of baselines in Recall@K, with the best score boldfaced and the second best underlined.
Base Methods CSFCube DORISMAE LitSearch
R@50 R@100 R@50 R@100 R@5 R@20 R@100
SPECTER-v2 Retriever 0.5331 0.6860 0.5305 0.7208 0.3931 0.5551 0.7205
BERT-QE 0.5243 0.6689 0.5284 0.6942 0.3959 0.5551 0.7364
ToTER 0.5443 0.7131 0.5319 0.7234 0.3948 0.5568 0.7239
HyDE 0.5879 0.7473 0.5110 0.6789 0.4241 0.5923 0.7682
GRF 0.5599 0.6758 0.5442 0.7283 0.4319 0.5830 0.7604
CSQE 0.5586 0.7149 0.5022 0.6491 0.4366 0.5747 0.7223
SemRank 0.6222 0.7601 0.5894 0.7451 0.5028 0.6316 0.7746
E5-large-v2 Retriever 0.6111 0.7362 0.5548 0.7162 0.5137 0.6573 0.7765
BERT-QE 0.6399 0.7589 0.5943 0.7451 0.4906 0.6361 0.7881
ToTER 0.6134 0.7553 0.5596 0.7120 0.4981 0.6627 0.7951
HyDE 0.6203 0.7255 0.5559 0.7236 0.4854 0.6592 0.8149
GRF 0.6183 0.7797 0.6025 0.7501 0.5347 0.6984 0.8319
CSQE 0.6416 0.7549 0.4787 0.5977 0.5503 0.6389 0.7719
SemRank 0.6661 0.8177 0.6286 0.7754 0.5807 0.7042 0.8312
Table 3: Performance of SemRank on three datasets with different base retrievers, with the best score boldfaced.
Methods CSFCube DORISMAE LitSearch
R@50 R@100 R@50 R@100 R@5 R@20 R@100
SPECTER-v2 0.5331 0.6860 0.5305 0.7208 0.3931 0.5551 0.7205
   + SemRank 0.6222 0.7601 0.5894 0.7451 0.5028 0.6316 0.7746
BM25 0.4651 0.5966 0.5721 0.7441 0.4381 0.5794 0.7362
   + SemRank 0.5840 0.7076 0.6507 0.8104 0.5126 0.6612 0.7920
Hybrid 0.5855 0.7131 0.6743 0.8297 0.5397 0.6877 0.7881
   + SemRank 0.6434 0.7673 0.6779 0.8358 0.5835 0.7161 0.7922
E5-large-v2 0.6111 0.7362 0.5548 0.7162 0.5137 0.6573 0.7765
   + SemRank 0.6661 0.8177 0.6286 0.7754 0.5807 0.7042 0.8312
GritLM-7B 0.6732 0.7742 0.6415 0.8037 0.6908 0.8001 0.9149
   + SemRank 0.7290 0.8466 0.6743 0.8187 0.6955 0.8171 0.9221
Base Retrievers
We use a wide range of base retrievers in our experiments. We include a sparse retriever BM25, an unsupervised dense retriever SPECTER-v2 Singh et al. (2022), and the Hybrid of these two. We also include two instruction-tuned dense retrievers, E5-large-v2 Wang et al. (2022) and GritLM-7B Muennighoff et al. (2025).
Baselines
We compare SemRank with a collection of retrieval methods using corpus knowledge and/or LLMs to enhance base retrievers.
  • BERT-QE Zheng et al. (2020) expands the query with relevant text chunks selected from top-ranked papers returned by the base retriever.
  • ToTER Kang et al. (2024a) improves the retrieval performance with a topical taxonomy by comparing the topic distributions of the query and documents predicted by a text classifier.
  • HyDE Gao et al. (2023) prompts an LLM to generate hypothetical document that answers the query and encode it as the query vector.
  • GRF Mackie et al. (2023) generates relevant context by an LLM. We choose to generate scientific concepts for fair comparison.
  • CSQE Lei et al. (2024) uses an LLM to select relevant sentences from top-ranked documents of the base retriever, which are then used to expand the query together with a hypothetical document.
Evaluation Metrics
We use Recall@K (R@K) as our evaluation metric. Following previous studies, we use $K=50,100$ for CSFCube and DORISMAE, and $K=5,20,100$ for LitSearch.
Implementation Details
For building semantic index, we train our text classifier using MAPLE Zhang et al. (2023b) which contains topic labels from Microsoft Academic Graph Sinha et al. (2015). We initialize it with SPECTER-v2-base and fine-tune it with learning rate at 5e-5 for 10 epochs. The balancing factor $\alpha$ is set to 1e-2. We also use SPECTER-v2-base for concept encoding. During retrieval, the number of candidate topics and key phrases is $k=50$. We use GPT-4.1-mini as the LLM for SemRank and all LLM-based baselines for fair comparison. The experiments are run on one NVIDIA RTX A6000 when a GPU is needed.  [^4] Code is available upon request and will later be public.
4.2 Retrieval Performance Comparison
Table 2 shows the results of compared baselines with two retrievers, SPECTER-v2 and E5-large. We clearly see that SemRank overall outperforms the compared baselines on all datasets. Specifically, we observe that methods using LLMs for query understanding achieves better performance, showing the strength of text understanding ability of LLMs in the retrieval task. Comparing with previous methods, SemRank uses the LLM for concept-based query understanding and matching, which captures the query’s need more explicitly and thus achieves stronger performance.
Table 3 additionally shows the results of SemRank on different base retrievers. We can see that SemRank consistently improves the performance of all kinds of retrievers. Even for GritLM-7B, the SOTA model reported in Ajith et al. (2024) and trained with scientific data, SemRank still improves its performance on all datasets. Besides, SemRank also improves the performance of Hybrid model, showing that concept-level semantic matching is not a combination of typical sparse and dense features, but in another intermediate level that really captures scientific knowledge.
4.3 Ablation Studies
We conduct ablation studies to show the effectiveness of each component of SemRank. We include the following ablated versions:
  • No Topic: excludes topics in semantic index.
  • No Phrase: excludes phrases in semantic index.
  • No Corpus: excludes the candidate scientific concepts from the corpus and thus prompts the LLM to directly generate scientific concepts based on its own knowledge.
  • No LLM (class): excludes the LLM and selects scientific concepts for the query using the fine-tuned topic classifier.
  • No LLM (freq): excludes the LLM and uses the top-20 most frequent concepts mentioned by top-ranked papers returned by the base retriever.
Table 4 shows the results of ablation studies on the LitSearch dataset with SPECTER-v2 as the base retriever. First, we observe that either ablating topics or phrases from the semantic index will degrade the performance, with phrases affecting more because of its capturing more detailed information. Second, removing the augmented corpus knowledge from LLM prompting will greatly affect the final performance, because LLMs tend to generate terms not matched with the corpus. Finally, removing LLM but using topic classifier or statistic-based metric for query concept identification also drastically decreases the performance, showing the power of LLM on query understanding when augmented with corpus knowledge.
Table 4: Ablation studies of SemRank on LitSearch.
R@5 R@20 R@100
SPECTER-v2 0.3931 0.5551 0.7205
Indexing
   No Topic 0.4331 0.6040 0.7687
   No Phrase 0.4160 0.5682 0.7447
Retrieval
   No Corpus 0.4060 0.5557 0.7260
   No LLM (class) 0.4079 0.5459 0.7320
   No LLM (freq) 0.3897 0.5497 0.7267
SemRank 0.5028 0.6316 0.7746
4.4 Efficiency Analysis
We show the efficiency of SemRank by comparing it with other LLM-based baselines. Specifically, we report the following factors of each method: the number of base retriever calls per query (# RET), the number of LLM calls per query (# LLM), the average number of tokens generated by the LLM per query (LLM Output Len), and the average running time per query (Running Time). As stated in Sect. 4.1, we use the same LLM checkpoint for all baselines for a fair comparison. Table 5 shows the detailed comparison results on LitSearch. We can clearly see that SemRank takes the least inference time among the compared methods, with 1.5$\times$ faster than the second fast method. Not only does SemRank call the retriever and LLM only once, it also expects minimal number of tokens responded by the LLM because of the concept-based structured input and output format.
4.5 Combination with LLM-based Reranking
Recent studies also use LLM for reranking retrieval results by prompting it to provide a new ranked list of top documents. To show that SemRank can naturally integrate with such a method, we compare the performance of LLM-based reranking with and without SemRank. Following Ajith et al. (2024), we provide top-100 papers to the LLM and use the prompt from Sun et al. (2023). Table 6 shows the results on LitSearch with 3 base retrievers. We can see SemRank does not conflict with LLM-based reranking by consistently improving its performance. While typical reranking only improves recall within provided number of papers (R@100 unchanged), SemRank can brings more relevant papers and also improves R@100.
Table 5: Efficiency analysis of LLM-based methods.
# RET # LLM LLM Output Len Running Time
HyDE 1 1 169.39 tok   4.02 sec
GRF 2 1   79.72 tok   2.78 sec
CSQE 2 2 462.21 tok 11.27 sec
SemRank 1 1    18.92 tok    1.82 sec
4.6 Parameter Studies
We study the influence of setting different number of candidates provided to LLM for query concept identification, i.e., $k$ in Sect. 3.2 for prompt in Figure 5. We set the value of $k=5,10,25,50,75,100$ and report their performance on LitSearch with base retriever SPECTER-v2 and GritLM. Results in Figure 3 show that SemRank overall is not very sensitive to the value of $k$ for $k\geq 25$. We notice slightly increased performance for smaller $k$ in the R@5 measures, which shows that an incomplete query concept set may not be sufficient to affect top-ranked documents that are hard to distinguish.
5 Related Works
Dense Retrieval
Dense retrieval has become a core paradigm in modern information retrieval. Early models like Dense Passage Retrieval (DPR) Karpukhin et al. (2020) and ME-BERT Luan et al. (2021) leveraged in-batch and BM25-based hard negatives to improve training efficiency. Subsequent methods refined negative sampling: ANCE Yu et al. (2021) used asynchronously updated indices; RocketQA Qu et al. (2021) introduced cross-batch and denoised negatives; and ADORE Zhan et al. (2021) adopted dynamic sampling for greater stability. PAIR Ren et al. (2021) incorporated passage-level similarity signals, while FiD-KD Izacard and Grave (2021) distilled knowledge from reader models.
In academic domains, specialized models like SciBERT Beltagy et al. (2019) pre-trained on scientific texts laid the groundwork. Recent innovations also focus on knowledge distillation from cross-encoder rankers  Huang and Chen (2024); Tao et al. (2024); Zhang et al. (2023a). Despite the progress and varied strategies for improving dense retrieval, these methods inherently rely on representing entire documents or queries with single dense embeddings, restraining them to capture fine-grained details crucial for accurately interpreting complex scientific queries. In contrast, our approach explicitly models the multi-granular scientific concepts within both queries and documents.
Table 6: Further analysis of SemRank by combining with LLM-based reranking.
R@5 R@20 R@100
SPECTER-v2 0.3931 0.5551 0.7205
   + Reranking 0.6636 0.7038 0.7205
    + SemRank 0.6705 0.7435 0.7746
E5-large-v2 0.5137 0.6573 0.7765
   + Reranking 0.6989 0.7480 0.7765
    + SemRank 0.7108 0.7963 0.8312
GritLM-7B 0.6908 0.8001 0.9149
   + Reranking 0.7575 0.8470 0.9149
    + SemRank 0.7774 0.8520 0.9221
LLM-Enhanced Retrieval
Recent works show that LLMs can boost retrieval quality even when little or no human supervision is available. In zero-shot settings, HyDE Gao et al. (2023) prompts an instruction-tuned LLM (e.g., InstructGPT) to imagine a hypothetical answer document, then encodes this synthetic text with an unsupervised dual-encoder; the dense representation guides nearest-neighbor search and already outperforms Contriever Izacard et al. (2021) without any task data. At inference time, LLMs refine the query itself. Rewrite-Retrieve-Read pipeline Ma et al. (2023) trains a lightweight rewriter with RL from the downstream reader LLM, consistently topping standard retrieve-then-read QA. For expansion, HyDE’s hypothetical-document trick underpins GenRead Yu et al. (2023), which sometimes matches or beats retrieval-based pipelines by generating context first. CSQE Lei et al. (2024) tempers hallucinations by mixing LLM expansions with sentences extracted from top-ranked corpus hits, outperforming fine-tuned neural expanders on tough TREC queries. CCQGen Kang et al. (2025) leverages an LLM to select topical concepts derived from a predefined top-down taxonomy. In our work, SemRank leverages an LLM to select core concepts directly from candidate sets of both broad research topics and specific key phrases derived from the corpus, which augment LLMs with multi-granular knowledge to help query understanding and reduce hallucination.
Retrieval with Corpus Knowledge
Recent work has explored leveraging corpus-based knowledge to enhance retrieval accuracy through query expansion and refinement techniques. Methods such as BERT-QE Zheng et al. (2020) select corpus-contextualized text chunks to alleviate vocabulary mismatches, while GAR Mao et al. (2021) generates pseudo-passages for semantic enrichment of queries, demonstrating substantial gains. Query2doc Wang et al. (2023b) further utilizes LLMs to create entire pseudo-documents, capturing external knowledge from web-scale training corpora. Similarly, classical pseudo-relevance feedback (PRF) Wang et al. (2021); Lei et al. (2024) has been adapted to dense retrieval, reducing hallucinations and improving effectiveness. Graph-based methods such as MacAvaney et al. (2022); Kulkarni et al. (2023) utilize document similarity graphs to dynamically expand search results. ToTER Kang et al. (2024a) designs taxonomy-based retrieval to identify the central topic classes and exploit their topical relatedness to supplement PLM-based retrievers. TaxoIndex Kang et al. (2024b) constructs a semantic index guided by an academic taxonomy, extracting and organizing concepts from documents, and then trains an indexing module to match these concepts with queries. In contrast, our proposed SemRank framework avoids supervision, utilizing corpus-derived concepts in an unsupervised manner to build semantic index and perform pseudo-relevance feedback, thereby enhancing retrieval without additional training.
Figure 3: Parameter analysis on LitSearch by varying $k$, the number of candidate query concepts.
6 Conclustion
We present SemRank, a novel scientific paper retrieval method that integrates LLM-guided query understanding with a concept-based semantic index. To overcome the limitations of existing methods, SemRank identifies multi-granular scientific concepts to explicitly understand scientific queries at the concept level. By augmenting LLMs with corpus knowledge, SemRank also facilitates LLM’s understanding of query and context while reducing hallucination. Experiments demonstrate that SemRank consistently improves retrieval performance across various base retrievers and outperforms various baseline methods while remaining highly efficient.
Limitations
While SemRank shows strong performance and efficiency in scientific paper retrieval, it also has several limitations. First, our current studies limit to scientific paper retrieval dataset with only title and abstract, while retrieving full scientific papers could be more challenging due to the difficulties of effectively understanding long structured text. Second, SemRank only consider scientific concepts as a set, while not considering their internal relationship which could bring more insights to paper and query understanding. Third, although our use of LLMs is efficient, the reliance on prompting still introduces sensitivity to prompt design and model behavior, which may require tuning for different domains. Finally, we only focus on English scientific paper retrieval in the work, while it remains a challenge on multilingual or multi-modal (e.g., figures, tables) paper retrieval.
Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329.